### v2

# Sharpness-Aware Minimization with Z-Score Gradient Filtering for Neural Networks

Abstract
Generalizing well in deep neural networks remains a core challenge, particularly due to their tendency to converge to sharp minima that degrade robustness. Sharpness-Aware Minimization (SAM) mitigates this by seeking flatter minima but perturbs parameters using the full gradient, which can include statistically insignificant directions. We propose ZSharp, a simple yet effective extension to SAM that applies layer-wise Z-score normalization followed by percentile-based filtering to retain only statistically significant gradient components. This selective perturbation aligns updates with curvature-sensitive directions, enhancing generalization without requiring architectural changes. ZSharp introduces only one additional hyperparameter, the percentile threshold, and remains fully compatible with existing SAM variants. Experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet using ResNet, VGG, and Vision Transformers show that ZSharp consistently outperforms SAM and its variants in test accuracy, particularly on deeper and transformer-based models. These results demonstrate that ZSharp is a principled and lightweight improvement for sharpness-aware optimization.
footnotetext:
1 Introduction
Deep neural networks (DNNs) [7] have demonstrated strong performance across a broad spectrum of machine learning tasks, including image classification [15], speech recognition [9], and natural language understanding [24]. These models are trained by minimizing an empirical loss over high-dimensional input data using gradient-based optimizers such as stochastic gradient descent (SGD) [22, 3]. While highly expressive, DNNs remain prone to overfitting, and improving their generalization to unseen data is a central challenge in deep learning research [30]. Several studies have linked poor generalization performance to convergence toward sharp minima—regions in the loss landscape characterized by high curvature [10, 12, 6]. In such regions, small perturbations to inputs or parameters may cause large increases in loss, suggesting reduced robustness [20]. This issue is particularly pronounced in overparameterized models, where the optimization trajectory has many possible directions, not all equally favorable for generalization [16].
To mitigate these effects, Sharpness-Aware Minimization (SAM) [6] was introduced, which encourages convergence to flatter minima by optimizing the worst-case loss within a neighborhood around the current parameters. SAM perturbs model weights in the direction of the gradient and updates them based on the maximum loss within a fixed-radius $\ell_{2}$ ball. This method improves generalization in various scenarios, particularly under large-batch training. However, a limitation of SAM is that it uses the entire gradient vector to determine the perturbation direction, including components with high variance or low statistical relevance. These components may contribute less to curvature-related features of the loss surface and can dilute the sharpness-sensitive signal [31].
In this work, we introduce Z-Score Filtered Sharpness-Aware Minimization (ZSharp), a simple yet effective refinement of SAM that incorporates statistical filtering into the sharpness-aware optimization process. ZSharp first applies layer-wise Z-score normalization [28] to standardize gradient distributions and then filters components based on a user-defined percentile threshold. By retaining only the top-ranked components (e.g., top 5%) based on their absolute Z-scores, the method targets gradient directions that exhibit high deviation relative to their layer-wise mean. These directions are assumed to be more informative for sharpness and curvature, aligning the perturbation with more geometrically relevant features in the loss surface.
Figure 1: Top-1 Test Accuracy on CIFAR-10 [14] with ResNet-56 [8] and ViT-7/8/8-384 [5]. ZSharp outperforms baseline AdamW [13, 19], SAM [6], Friendly-SAM [18], and ASAM [16], demonstrating improved generalization across architectures.
Unlike prior variants such as ASAM [16], which adjusts the perturbation magnitude based on estimated local curvature, or Friendly-SAM [18], which approximates the sharpness-aware objective to reduce computation, ZSharp directly refines the perturbation direction through statistically grounded sparsification. Importantly, ZSharp introduces only one additional hyperparameter-the percentile threshold-on top of SAM’s perturbation radius. It does not require changes to the model architecture or the underlying optimizer, making it easy to integrate into existing training pipelines. As shown in Figure 1, ZSharp achieves higher top-1 test accuracy compared to SAM [6], ASAM [16], and Friendly-SAM [18], when evaluated on CIFAR-10 [14] using both ResNet-56 and ViT [5] architectures. These gains are observed across architectures, including transformer-based models such as ViT, where gradient noise is typically more prevalent due to high-dimensional parameter spaces. In such settings, filtering noisy gradients may lead to more stable and effective updates.
To summarize, ZSharp offers a statistically motivated extension of SAM that enhances sharpness-aware perturbation through simple yet principled gradient filtering. Its sparsification mechanism improves signal quality without significantly increasing computational overhead or tuning complexity. Below, we highlight the key contributions of this work:
Our contributions are summarized as follows:
  • We propose ZSharp, a sharpness-aware optimization method that filters gradient components based on Z-score normalization to emphasize directions with high statistical significance.
  • ZSharp introduces only a single percentile threshold as an additional hyperparameter and retains the simplicity and compatibility of the original SAM framework.
To evaluate ZSharp, we conduct experiments on three widely used image classification benchmarks-CIFAR-10 [14], CIFAR-100 [14], and Tiny-ImageNet [17]-covering tasks of varying difficulty and granularity. We apply ZSharp to multiple architectures, including convolutional networks (ResNet [8], VGG [23]) and Vision Transformers (ViT) [5]. Across all settings, ZSharp demonstrates improved generalization over baseline SAM and its variants. These results suggest that statistically guided filtering improves the sharpness-sensitivity of perturbations, particularly in noisy or high-dimensional optimization regimes.
2 Related Works
The ongoing effort to improve generalization in deep neural networks (DNNs) [7] has motivated the development of numerous optimization strategies, particularly those that account for the geometry of the loss landscape. Among these, sharpness-aware optimization has emerged as a prominent framework, aiming to identify flatter minima that are empirically associated with improved generalization. In this section, we review representative approaches in sharpness-aware optimization and gradient normalization, and position our proposed method, ZSharp, in the context of this prior work.
Sharpness-Aware Minimization (SAM) [6] is a widely recognized method in this domain. It promotes flatter minima by optimizing the worst-case loss within a local neighborhood around the model parameters. Specifically, SAM perturbs the parameters in the direction of the gradient and minimizes the maximum loss within an $\ell_{2}$-ball. This approach has demonstrated improved generalization over standard optimizers such as Adam in several image classification benchmarks [4]. However, SAM constructs its perturbation using the entire gradient vector, which includes components that may be less informative or statistically noisy in the context of curvature. This can reduce the precision of perturbations when identifying sharpness-sensitive directions [31].
To address these concerns, a number of SAM variants have been introduced. Adaptive SAM (ASAM) [16] adjusts the perturbation scale according to the curvature of the loss surface, increasing robustness to parameter scaling. Friendly-SAM [18] reduces the computational cost of SAM by approximating the sharpness objective using a projected gradient. While this approximation improves efficiency, it may trade off some accuracy, particularly in architectures with complex parameter interactions such as Vision Transformers (ViTs) [5]. GSAM [31] introduces gradient alignment to stabilize sharpness-aware updates but adds hyperparameters that may increase tuning complexity.
Beyond sharpness-aware methods, several normalization techniques have been proposed to improve training stability and generalization. Activation-based approaches such as BatchNorm [11] and LayerNorm [1] normalize intermediate representations to reduce internal covariate shift. In contrast, gradient-level normalization directly adjusts the update step. Gradient clipping [21] constrains gradient magnitude to prevent instability, while gradient centralization [25] improves convergence by subtracting the mean from gradients. Stochastic Gradient Sampling, as introduced in StochGradAdam [29, 27], selectively samples a subset of gradients during training to enhance computational efficiency while maintaining performance comparable to Adam. More recently, ZNorm [28, 26] applies layer-wise Z-score normalization to gradients, ensuring consistent scaling across layers and showing promising results on datasets like CIFAR-10 and in medical imaging applications.
Figure 2: Overview of ZSharp: Z-Score Filtered Sharpness-Aware Minimization. (A) Gradients from fully-connected and convolutional layers are used to compute layer-wise statistics. (B) Z-score normalization is applied to standardize gradients, followed by percentile-based filtering to select statistically significant components. (C) A binary mask retains only the top Z-score entries (e.g., top 5%), filtering the gradient for the ascent step. (D) The filtered gradient is then used in the SAM ascent phase to refine the perturbation direction, enhancing generalization by focusing updates on curvature-sensitive directions.
ZSharp builds on the foundations of SAM [6] and ZNorm [28, 26] by introducing statistical filtering into the perturbation step. During the ascent phase, ZSharp applies layer-wise Z-score normalization to the gradient and filters out components below a user-defined percentile threshold. By retaining only the top $(1-Q_{p})\%$ of components with high standardized values, which $Q_{p}$ means percentile, ZSharp emphasizes gradient directions that deviate substantially from the layer mean—directions that may carry stronger geometric relevance. This filtering yields sparse but targeted perturbations that aim to better capture sharpness-related directions in the loss landscape.
Importantly, ZSharp introduces only one additional hyperparameter—the percentile threshold—and does not require modifications to the base architecture or training objective. It remains compatible with existing SAM implementations and base optimizers. Empirical results on image classification benchmarks including CIFAR-10 [14], CIFAR-100 [14], and Tiny-ImageNet [17] suggest that ZSharp achieves improved generalization across various architectures, including convolutional networks like ResNet [8] and VGG [23], as well as transformer-based models such as ViT [5]. These findings support the view that statistically guided filtering may be an effective strategy to enhance sharpness-aware optimization, particularly in high-dimensional or noisy gradient regimes.
3 Methodology
We propose Z-Score Filtered Sharpness-Aware Minimization (ZSharp), a novel method that enhances neural network training by integrating Z-score Normalization into a sharpness-aware framework. ZSharp filters gradients to focus on statistically significant directions (gradient directions with high magnitude), optimizing for flatter minima to improve generalization. An overview of our ZSharp method is provided in Figure 2. Below, we detail Z-score gradient normalization [28] and the Z-score filtered sharpness-aware minimization process.
Preliminaries. We consider a supervised learning framework with a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{m}$ denotes input features and $y_{i}\in\mathcal{Y}$ represents labels. The neural network, parameterized by weights $\theta\in\mathbb{R}^{d}$, defines a mapping $f:\mathbb{R}^{m}\times\mathbb{R}^{d}\to\mathcal{Y}$. For $L$ layers, the $\ell$-th layer’s parameters are $\theta^{(\ell)}\in\mathbb{R}^{d_{\ell}}$, with $\sum_{\ell=1}^{L}d_{\ell}=d$. The empirical loss $\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\ell(f(\mathbf{x}_{i};\theta),y_{i})$ has gradient $\nabla\mathcal{L}(\theta)\in\mathbb{R}^{d}$, with $\ell_{2}$-norm $\|\nabla\mathcal{L}(\theta)\|_{2}$. Sharp minima often hinder generalization [10], motivating our ZSharp method to favor flatter minima by filtering gradients based on a percentile threshold $Q_{p}$, retaining $(1-Q_{p})$ of the original gradients with the largest magnitudes.
Layer-wise Z-Score Gradient Normalization. Z-score Gradient Normalization (ZNorm) [28] standardizes gradient statistics for each layer $\ell$ independently, preserving the network’s layer-wise architecture. The gradient vector for the $\ell$-th layer is denoted $\nabla\mathcal{L}^{(\ell)}(\theta)\in\mathbb{R}^{d_{\ell}}$. This process is illustrated in (A) in Figure 2. The layer-wise gradient mean and standard deviation are computed as:
$\mu^{(\ell)}$ $=\frac{1}{d_{\ell}}\sum_{j=1}^{d_{\ell}}{\nabla\mathcal{L}^{(\ell)}(\theta)}_{j},$ (1)
$\sigma^{(\ell)}$ $=\left(\frac{1}{d_{\ell}}\sum_{j=1}^{d_{\ell}}\left({\nabla\mathcal{L}^{(\ell)}(\theta)}_{j}-\mu^{(\ell)}\right)^{2}\right)^{1/2}$ (2)
providing a normalized representation of gradient components per layer. The Z-score normalized gradient ${\Omega}({\nabla\mathcal{L}^{(\ell)}(\theta)})\in\mathbb{R}^{d_{\ell}}$ is defined as:
$\Omega({\nabla\mathcal{L}^{(\ell)}(\theta)})_{j}=\frac{{\nabla\mathcal{L}^{(\ell)}(\theta)}_{j}-\mu^{(\ell)}}{\sigma^{(\ell)}+\delta},\quad\forall j\in\{1,2,\dots,d_{\ell}\},$ (3)
where $\delta=10^{-8}$ ensures numerical stability. In vector form:
$\mathbf{\Omega}^{(\ell)}=\frac{\mathbf{\nabla\mathcal{L}^{(\ell)}(\theta)}-\mu^{(\ell)}\cdot\mathbf{1}}{\sigma^{(\ell)}+\delta},$ (4)
with $\mathbf{1}\in\mathbb{R}^{d_{\ell}}$ representing the all-ones vector. This normalization per layer ensures $\mathbb{E}[\mathbf{\Omega}^{(\ell)}]\approx 0$ and $\text{Var}[\mathbf{\Omega}^{(\ell)}]\approx 1$, enabling the detection of statistically significant gradients within each layer. Z-score normalization enables percentile-based filtering to identify statistically significant gradients by standardizing their distribution, whereas direct percentile filtering of raw gradients could be biased by layer-specific scale variations.
3.1 Z-Score Filtered Sharpness-Aware Minimization
Z-Score Filtered Sharpness-Aware Minimization (ZSharp) combines ZNorm [28] with sharpness-aware minimization [6] to refine gradient perturbations, explicitly targeting flatter regions of the loss landscape. After applying ZNorm, ZSharp filters gradients based on a percentile threshold, retaining $(1-Q_{p})$ of the most significant components; for example, with $Q_{p}=0.95$, the top 5% of gradients with the largest magnitudes are kept for the ascent step. This normalization and filtering process corresponds to steps (B) and (C) in the ZSharp overview (Figure 2). The method operates in two phases: an ascent step with ZSharp to assess sharpness and a descent step to minimize the loss.
Sharpness Awareness and Ascent. ZSharp minimizes the sharpness of the loss landscape, defined as the worst-case loss within a local $\ell_{2}$-ball of radius $\rho>0$ around the parameters [6]:
$\min_{\theta\in\mathbb{R}^{d}}\sup_{\epsilon\in\mathcal{B}_{2}(\rho)}\mathcal{L}(\theta+\epsilon),$ (5)
where $\mathcal{B}_{2}(\rho)=\{\epsilon\in\mathbb{R}^{d}\mid\|\epsilon\|_{2}\leq\rho\}$ is the local perturbation ball. ZSharp refines the perturbation direction by filtering $\nabla\mathcal{L}(\theta)$ via Z-score normalization, retaining only the gradients with the highest Z-scores.
Let $\Omega(\nabla\mathcal{L}(\theta))\in\mathbb{R}^{d}$ be the Z-score normalized gradient. A binary mask $\mathbf{m}\in\{0,1\}^{d}$ is applied, where:
$m_{j}=\begin{cases}1&\text{if }|\Omega(\nabla\mathcal{L}(\theta))_{j}|>q_{Q_{p}},\\ 0&\text{otherwise},\end{cases}\quad\forall j\in\{1,2,\dots,d\},$ (6)
where $q_{Q_{p}}$ is the Z-score value at the $Q_{p}$-th percentile of the absolute Z-scores. In other words, $q_{Q_{p}}$ represents the threshold Z-score corresponding to the $Q_{p}$-th percentile in the distribution of $|\Omega(\nabla\mathcal{L}(\theta))|$, ensuring that only the top $(1-Q_{p})$ fraction of gradients are retained. For example, in this work, we set $Q_{p}=0.95$, so $q_{Q_{p}}$ is the Z-score value at the 95th percentile, retaining the top $5\%$ of components with the largest Z-scores. This distribution is illustrated in Subfigure (B) of Figure 2, where red bars represent the top $5\%$, orange and red the top $10\%$, and yellow, orange, and red the top $15\%$, while gray bars correspond to gradients filtered out. The filtered gradient is (here, $\nabla\mathcal{L}(\theta)$ is the original gradient):
$\nabla\mathcal{L}(\theta)_{\text{filtered}}=\nabla\mathcal{L}(\theta)\odot\mathbf{m}.$ (7)
Following the approach in SAM [6], ZSharp approximates sharpness using the filtered gradient in a Taylor expansion:
$S(\theta)\approx\sup_{\|\epsilon\|_{2}\leq\rho}\left[\mathcal{L}(\theta)+\langle\nabla\mathcal{L}(\theta)_{\text{filtered}},\epsilon\rangle+\frac{1}{2}\epsilon^{\top}\mathcal{H}(\theta)\epsilon\right],$ (8)
where $\mathcal{H}(\theta)=\nabla^{2}\mathcal{L}(\theta)$ is the Hessian matrix, capturing the curvature of the loss landscape. The term $\langle\nabla\mathcal{L}(\theta)_{\text{filtered}},\epsilon\rangle$ accounts for the linear change in the loss along the filtered gradient direction, while $\frac{1}{2}\epsilon^{\top}\mathcal{H}(\theta)\epsilon$ quantifies the curvature, amplifying the influence of directions with high curvature (corresponding to large eigenvalues of $\mathcal{H}(\theta)$), which contribute significantly to sharpness. Unlike SAM, which uses the full gradient $\nabla\mathcal{L}(\theta)$, ZSharp leverages the filtered gradient $\nabla\mathcal{L}(\theta)_{\text{filtered}}$ to focus on directions with high Z-scores, ensuring that the perturbation $\epsilon$ aligns with significant components while suppressing noisy or insignificant ones. In practice, the Hessian is not computed directly due to its computational cost; instead, its effect is approximated through techniques like Hessian-vector products, focusing on curvature along the perturbation direction $\epsilon$.
The perturbation $\epsilon\in\mathbb{R}^{d}$, adapted from SAM [6] but using the filtered gradient, is defined as:
$\epsilon=\begin{cases}\rho\cdot\frac{\nabla\mathcal{L}(\theta)_{\text{filtered}}}{\|\nabla\mathcal{L}(\theta)_{\text{filtered}}\|_{2}+\delta}&\text{if }\|\nabla\mathcal{L}(\theta)_{\text{filtered}}\|_{2}>0,\\ \rho\cdot\frac{\nabla\mathcal{L}(\theta)}{\|\nabla\mathcal{L}(\theta)\|_{2}+\delta}&\text{otherwise},\end{cases}$ (9)
where $\delta=10^{-8}$ ensures numerical stability by preventing division by zero. This perturbation $\epsilon$ is a scaled version of the filtered gradient, aligned with significant directions identified by Z-score filtering, and has a magnitude controlled by the radius $\rho$.
The updated parameter becomes:
$\tilde{\theta}=\theta+\epsilon.$ (10)
This ascent step, inherited from SAM, guides optimization toward flatter minima by increasing the loss in directions of high sharpness, but ZSharp enhances this by focusing only on sharpness-critical directions with high Z-scores. The perturbed loss is then approximated as:
$\mathcal{L}(\tilde{\theta})\approx\mathcal{L}(\theta)+\langle\nabla\mathcal{L}(\theta)_{\text{filtered}},\epsilon\rangle+\frac{1}{2}\epsilon^{\top}\mathcal{H}(\theta)\epsilon,$ (11)
reflecting the combined effect of the filtered gradient and curvature along the perturbation direction.
Minimization and Descent. The descent phase of ZSharp minimizes the loss $\mathcal{L}:\mathbb{R}^{d}\to\mathbb{R}$ by leveraging the gradient at the perturbed parameters $\tilde{\theta}\in\mathbb{R}^{d}$. We compute the perturbed gradient:
$\nabla\mathcal{L}(\tilde{\theta})\in\mathbb{R}^{d},$ (12)
where $\tilde{\theta}=\theta+\epsilon$ is derived from the ascent step, and $\nabla$ denotes the gradient operator in $\mathbb{R}^{d}$. The original parameters $\theta^{(t)}\in\mathbb{R}^{d}$ at iteration $t\in\mathbb{N}$ are updated via a base optimizer $\mathcal{O}$, such as stochastic gradient descent (SGD) or Adam, defined as:
$\theta^{(t+1)}=\theta^{(t)}-\eta\mathcal{O}(\nabla\mathcal{L}(\tilde{\theta})),$ (13)
where $\eta>0$ is the learning rate, and $\mathcal{O}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ encapsulates the optimization rule (e.g., $\mathcal{O}(\nabla\mathcal{L}(\tilde{\theta}))\text{ means }\nabla\mathcal{L}(\tilde{\theta})$ is updated with a base optimizer such as SGD or Adam). This descent step drives $\theta$ towards the global minimizer:
$\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{d}}\mathcal{L}(\theta),$ (14)
while the prior ascent constrains the sharpness, quantified by $\sup_{\|\epsilon\|_{2}\leq\rho}\mathcal{L}(\theta+\epsilon)$. The two-phase process—sharpness-aware ascent followed by descent—optimizes $\mathcal{L}(\theta)$ over $\mathbb{R}^{d}$, balancing loss reduction with curvature minimization, as reflected in the Hessian $\mathcal{H}(\theta)=\nabla^{2}\mathcal{L}(\theta)\in\mathbb{R}^{d\times d}$.
Figure 3: Comparison of SAM and ZSharp updates on the loss landscape. Starting from $\theta_{t}$, SAM moves to $\tilde{\theta}$ (blue dashed ascent arrow) and then to $\theta^{\text{SAM}}_{t+1}$ (blue descent arrow). ZSharp moves to $\hat{\theta}$ (red dashed ascent arrow) using filtered gradients, focusing on high Z-score directions, and reaches $\theta^{\text{ZSharp}}_{t+1}$ (red descent arrow), guiding optimization toward flatter minima more efficiently. Standard gradient descent ($-\eta\nabla L(\theta_{t})$) is shown as a black dashed arrow for reference.
Weight Update Dynamics. The weight update in ZSharp integrates ascent and descent dynamics into a composite operation over the parameter space $\mathbb{R}^{d}$. The update rule combines the perturbation $\epsilon\in\mathbb{R}^{d}$ with the descent step:
$\theta^{(t+1)}=\theta^{(t)}+\epsilon-\eta\nabla\mathcal{L}(\theta^{(t)}+\epsilon),$ (15)
where $\theta^{(t)}\in\mathbb{R}^{d}$ is the parameter vector at iteration $t\in\mathbb{N}$, $\eta>0$ is the learning rate, and $\nabla\mathcal{L}(\theta^{(t)}+\epsilon)\in\mathbb{R}^{d}$ is the gradient at the perturbed point.
Substituting $\boldsymbol{\epsilon}$ from the ascent phase, for the case where $\|\nabla\mathcal{L}(\theta^{(t)})_{\text{filtered}}\|_{2}>0$, we expand the update as:
$\theta^{(t+1)}$ $=\theta^{(t)}+\rho\cdot\frac{\nabla\mathcal{L}(\theta^{(t)})_{\text{filtered}}}{\|\nabla\mathcal{L}(\theta^{(t)})_{\text{filtered}}\|_{2}+\delta}$
$\quad-\eta\nabla\mathcal{L}\left(\theta^{(t)}+\rho\cdot\frac{\nabla\mathcal{L}(\theta^{(t)})_{\text{filtered}}}{\|\nabla\mathcal{L}(\theta^{(t)})_{\text{filtered}}\|_{2}+\delta}\right),$ (16)
where $\nabla\mathcal{L}(\theta^{(t)})_{\text{filtered}}=\nabla\mathcal{L}(\theta^{(t)})\odot\mathbf{m}\in\mathbb{R}^{d}$, $\mathbf{m}\in\{0,1\}^{d}$ is the Z-score filtering mask, $\rho>0$ is the perturbation radius, and $\delta=10^{-8}$ ensures numerical stability. Letting $\mathbf{p}=\nabla\mathcal{L}(\theta^{(t)})_{\text{filtered}}$, the ascent contribution becomes:
$\boldsymbol{\epsilon}=\rho\cdot\frac{\mathbf{p}}{\|\mathbf{p}\|_{2}+\delta},$ (17)
yielding the full update:
$\theta^{(t+1)}$ $=\theta^{(t)}+\boldsymbol{\epsilon}-\eta\nabla\mathcal{L}(\theta^{(t)}+\boldsymbol{\epsilon})$
$=\theta^{(t)}+\rho\cdot\frac{\mathbf{p}}{\|\mathbf{p}\|_{2}+\delta}-\eta\nabla\mathcal{L}\left(\theta^{(t)}+\rho\cdot\frac{\mathbf{p}}{\|\mathbf{p}\|_{2}+\delta}\right).$ (18)
In the extreme case where the percentile threshold is set to $Q_{p}=1.0$, the corresponding Z-score threshold $q_{Q_{p}}$ becomes the maximum absolute Z-score. As a result, no component satisfies $|\Omega(\nabla\mathcal{L}(\theta))_{j}|>q_{Q_{p}}$, and thus the binary mask becomes:
$m_{j}=\begin{cases}1&\text{if }|\Omega(\nabla\mathcal{L}(\theta))_{j}|>q_{Q_{p}},\\ 0&\text{otherwise}\end{cases}\quad\Rightarrow\quad\mathbf{m}=\mathbf{0}.$
Consequently, the filtered gradient vanishes, $\nabla\mathcal{L}(\theta)_{\text{filtered}}=\nabla\mathcal{L}(\theta)\odot\mathbf{m}=\mathbf{0}$, and the perturbation becomes zero, $\boldsymbol{\epsilon}=\mathbf{0}$. In this degenerate case, the ascent step has no effect, and the update reduces to a standard gradient descent step:
$\theta^{(t+1)}=\theta^{(t)}-\eta\nabla\mathcal{L}(\theta^{(t)}).$ (19)
In contrast, when the percentile threshold is set to $Q_{p}=0$, the corresponding Z-score threshold $q_{Q_{p}}$ becomes zero, and all components satisfy $|\Omega(\nabla\mathcal{L}(\theta))_{j}|>q_{Q_{p}}$. Therefore, the mask becomes $\mathbf{m}=\mathbf{1}$, and the filtered gradient reduces to the full gradient:
$\nabla\mathcal{L}(\theta)_{\text{filtered}}=\nabla\mathcal{L}(\theta).$
In this case, ZSharp recovers the original SAM update rule:
$\theta^{(t+1)}$ $=\theta^{(t)}+\rho\cdot\frac{\nabla\mathcal{L}(\theta^{(t)})}{\|\nabla\mathcal{L}(\theta^{(t)})\|_{2}+\delta}$
$\quad-\eta\nabla\mathcal{L}\left(\theta^{(t)}+\rho\cdot\frac{\nabla\mathcal{L}(\theta^{(t)})}{\|\nabla\mathcal{L}(\theta^{(t)})\|_{2}+\delta}\right).$ (20)
This update combines the sharpness-aware ascent step with the descent step based on the perturbed gradient, guiding the model toward flatter minima. As $Q_{p}$ increases, ZSharp adaptively filters out less significant gradient directions, providing a strict generalization of SAM where the perturbation is selectively focused on statistically significant components identified by Z-score normalization. However, when SAM uses the full gradient for the ascent step, the perturbation may align with directions of high sharpness, potentially leading the optimization away from flatter minima, as the full gradient includes noisy or less significant components that can amplify curvature in undesirable directions. In contrast, ZSharp mitigates this by focusing solely on components with high Z-scores, ensuring that the perturbation targets sharpness-critical directions while preserving alignment with flatter regions of the loss landscape. This behavior is illustrated in Figure 3, where ZSharp’s update path (red arrows) more effectively navigates toward flatter minima compared to SAM (blue arrows), demonstrating the advantage of Z-score-based gradient filtering in achieving better generalization.
4 Experimental Results
Figure 4: Top-1 Test Accuracy comparison on CIFAR-10 for ResNet-56/110 and ViT-7/8/8-384 models across different SAM variants: AdamW (Baseline) [19], SAM [6], Friendly-SAM [18], ASAM [16], and ZSharp (Ours). The red dashed line indicates the baseline performance using AdamW [19] alone, highlighting the improvements achieved by sharpness-aware methods. ZSharp consistently outperforms other methods, demonstrating the effectiveness of ZNorm-based gradient filtering in enhancing generalization.
To assess the efficacy of ZSharp, we conducted a series of experiments comparing its performance against Sharpness-Aware Minimization (SAM) [6] and SAM variants such as ASAM [16], Friendly-SAM [18] and baseline optimizers. Our evaluation focuses on generalization performance, measured through test accuracy and robustness, across diverse datasets and model architectures. The following subsections detail our experimental setup and results, providing the performances of our proposed method.
Experimental Settings. We evaluate ZSharp on three standard benchmarks: CIFAR-10 [14], CIFAR-100 [14], and Tiny-ImageNet [17]. CIFAR-10 and CIFAR-100 each consist of 50,000 training and 10,000 test images at 32$\times$32 resolution across 10 and 100 classes, respectively. Tiny-ImageNet contains 90,000 training and 10,000 test images of size 64$\times$64 across 200 classes. We benchmark ZSharp using ResNet-56/110 [8], VGG16_BN [23]. For ViT models, ViT-7/8/8-384 and ViT-7/8/12-768 denote Vision Transformers [5] with 7 layers, 8 attention heads, patch sizes of 8, and MLP dimensions of 384 and 768, respectively. All trained without pre-trained weights.
All models are trained for 200 epochs with a batch size of 256 using the AdamW optimizer [13, 19] (initial learning rate 0.001, weight decay $5\times 10^{-5}$). The learning rate is scheduled using a step decay policy, where it is multiplied by 0.75 every 10 epochs. For SAM [6], ASAM [16], Friendly-SAM [18] and ZSharp, the perturbation radius $\rho$ is set to 0.05 as used in each paper, and we followed optimal hyperparameters written in papers. ZSharp uses Z-score filtering with $Q_{p}=0.95$, retaining only the top 5% of gradient components during the ascent step. All experiments are conducted on a single NVIDIA RTX 4090 GPU, and results are averaged over 3 different random seeds to ensure statistical robustness.
4.1 Experimental Results
CIFAR-10 [14] CIFAR-100 [14] Tiny-ImageNet [17] Network Method Top-1 Test Acc. Train Loss. Top-1 Test Acc. Train Loss. Top-1 Test Acc. Train Loss. ResNet-56 [8] AdamW (Baseline) [19] 0.9108 ± 0.0045 0.0057 ± 0.0013 0.6420 ± 0.0031 0.0452 ± 0.0629 0.4747 ± 0.0031 0.0121 ± 0.0105 SAM [6] 0.9160 ± 0.0021 0.0221 ± 0.0051 0.6527 ± 0.0025 0.1097 ± 0.0584 0.4938 ± 0.0026 0.0453 ± 0.0127 ASAM [16] 0.9228 ± 0.0034 0.0366 ± 0.0093 0.6646 ± 0.0017 0.1952 ± 0.0419 0.5072 ± 0.0012 0.0564 ± 0.0091 Friendly-SAM [18] 0.9179 ± 0.0025 0.0219 ± 0.0076 0.6549 ± 0.0024 0.1051 ± 0.0417 0.4948 ± 0.0023 0.0444 ± 0.0102 ZSharp (Ours) 0.9264 ± 0.0032 0.0630 ± 0.0064 0.6679 ± 0.0015 0.2510 ± 0.0438 0.5073 ± 0.0014 0.0828 ± 0.0129 ResNet-110 [8] AdamW (Baseline) [19] 0.9140 ± 0.0031 0.0056 ± 0.0023 0.6650 ± 0.0025 0.0149 ± 0.0059 0.4878 ± 0.0045 0.0556 ± 0.0114 SAM [6] 0.9233 ± 0.0025 0.0188 ± 0.0037 0.6815 ± 0.0019 0.0531 ± 0.0121 0.5005 ± 0.0045 0.1417 ± 0.0216 ASAM [16] 0.9261 ± 0.0023 0.0288 ± 0.0056 0.6796 ± 0.0036 0.0915 ± 0.0123 0.5105 ± 0.0045 0.2894 ± 0.0241 Friendly-SAM [18] 0.9193 ± 0.0013 0.0190 ± 0.0036 0.6762 ± 0.0021 0.0524 ± 0.0113 0.5027 ± 0.0045 0.1402 ± 0.0091 ZSharp (Ours) 0.9293 ± 0.0017 0.0618 ± 0.0097 0.6844 ± 0.0023 0.1656 ± 0.0213 0.5207 ± 0.0045 0.4137 ± 0.0311 VGG-16/BN [23] AdamW (Baseline) [19] 0.9247 ± 0.0013 0.0058 ± 0.0031 0.6999 ± 0.0102 0.0092 ± 0.0051 0.5507 ± 0.0093 0.0043 ± 0.0071 SAM [6] 0.9337 ± 0.0018 0.0171 ± 0.0093 0.7092 ± 0.0093 0.0139 ± 0.0073 0.5587 ± 0.0103 0.0363 ± 0.0183 ASAM [16] 0.9355 ± 0.0012 0.0237 ± 0.0047 0.7170 ± 0.0121 0.0375 ± 0.0118 0.5647 ± 0.0191 0.0644 ± 0.0237 Friendly-SAM [18] 0.9290 ± 0.0017 0.0163 ± 0.0093 0.7099 ± 0.0083 0.0495 ± 0.0125 0.5544 ± 0.0204 0.0349 ± 0.0153 ZSharp (Ours) 0.9327 ± 0.0020 0.0351 ± 0.0144 0.7207 ± 0.0071 0.0375 ± 0.0137 0.5673 ± 0.0231 0.1248 ± 0.0351 ViT-7/8/8-384 [5] AdamW (Baseline) [19] 0.8398 ± 0.0028 0.0087 ± 0.0092 0.5479 ± 0.0011 0.0042 ± 0.0031 0.2843 ± 0.0012 0.0056 ± 0.0014 SAM [6] 0.8432 ± 0.0032 0.0273 ± 0.0101 0.5557 ± 0.0013 0.0255 ± 0.0141 0.2897 ± 0.0009 0.0363 ± 0.0098 ASAM [16] 0.8302 ± 0.0034 0.0367 ± 0.0138 0.5566 ± 0.0031 0.0349 ± 0.0193 0.2522 ± 0.0032 0.0644 ± 0.0137 Friendly-SAM [18] 0.8476 ± 0.0044 0.0273 ± 0.0093 0.5608 ± 0.0023 0.0228 ± 0.0138 $0$ 0.0012 0.0349 ± 0.0083 ZSharp (Ours) 0.8543 ± 0.0029 0.0647 ± 0.0216 0.5748 ± 0.0051 0.0730 ± 0.0212 0.3057 ± 0.0021 0.1248 ± 0.0413 ViT-7/8/12-768 [5] AdamW (Baseline) [19] 0.8438 ± 0.0021 0.0087 ± 0.0031 0.5615 ± 0.0013 0.0040 ± 0.0045 0.2991 ± 0.0010 0.0065 ± 0.0032 SAM [6] 0.8486 ± 0.0018 0.0293 ± 0.0098 0.5691 ± 0.0014 0.0234 ± 0.0076 0.3014 ± 0.0015 0.0297 ± 0.0098 ASAM [16] 0.8395 ± 0.0020 0.0371 ± 0.0101 0.5649 ± 0.0027 0.0347 ± 0.0116 0.3023 ± 0.0008 0.0512 ± 0.0161 Friendly-SAM [18] 0.8525 ± 0.0021 0.0283 ± 0.0084 0.5655 ± 0.0021 0.0246 ± 0.0122 0.3034 ± 0.0013 0.0441 ± 0.0141 ZSharp (Ours) 0.8586 ± 0.0023 0.0635 ± 0.0196 0.5777 ± 0.0031 0.0709 ± 0.0178 0.3104 ± 0.0019 0.1341 ± 0.0211
Table 1: Top-1 Test Accuracy and Train Loss for ResNet-56 [8], ResNet-110 [8], VGG-16/BN [23], ViT-7/8/8-384 [5], and ViT-7/8/12-768 [5] on CIFAR-10 [14], CIFAR-100 [14], and Tiny-ImageNet datasets [17] across different SAM variants such as AdamW (Baseline) [19], SAM [6], Friendly-SAM [18], ASAM [16], and ZSharp (Ours). For ViT models, ViT-7/8/8-384 and ViT-7/8/12-768 denote Vision Transformers with 7 layers, 8 attention heads, patch sizes of 8, and MLP dimensions of 384 and 768, respectively.
To evaluate ZSharp, we compare its performance, measured by test accuracy, against SAM [6], ASAM [16], Friendly-SAM [18], and AdamW (Baseline) [19] across diverse datasets and architectures. Table 1 reports the Top-1 Test Accuracy for ResNet-56, ResNet-110, VGG-16/BN, ViT-7/8/8-384, and ViT-7/8/12-768 on CIFAR-10, CIFAR-100, and Tiny-ImageNet. ZSharp achieves the highest test accuracy on ResNet-56 (CIFAR-10: 0.9264 ± 0.0032 vs. SAM’s 0.9160 ± 0.0021), ResNet-110 (CIFAR-100: 0.6844 ± 0.0023 vs. SAM’s 0.6815 ± 0.0019), VGG-16/BN (CIFAR-100: 0.7207 ± 0.0071 vs. SAM’s 0.7092 ± 0.0093), ViT-7/8/8-384 (Tiny-ImageNet: 0.3057 ± 0.0021 vs. SAM’s 0.2897 ± 0.0009), and ViT-7/8/12-768 (CIFAR-10: 0.8586 ± 0.0023 vs. SAM’s 0.8486 ± 0.0018).
Figure 4 shows the test accuracy curves over epochs for ResNet-56, ResNet-110, and ViT-7/8/8-384 on CIFAR-10, CIFAR-100, and Tiny-ImageNet, comparing ZSharp with SAM variants and the baseline.
4.2 Hyperparameter Tuning
We evaluate the effect of the percentile threshold $Q_{p}$ in ZSharp, which controls the proportion of gradient components retained after ZNorm filtering, selecting the top $(1-Q_{p})\%$ based on Z-scores. A higher $Q_{p}$ (e.g., 0.95) retains fewer components (top 5%), focusing on significant directions for sharpness-aware optimization, while a lower $Q_{p}$ (e.g., 0.75) retains more (top 25%). At $Q_{p}=0.0$, ZSharp reduces to SAM, using the full gradient. We test $Q_{p}\in\{0.75,0.80,0.85,0.90,0.95\}$ on ResNet-56 and ViT-7/8/8-384, with results in Table 2.
Network Method $Q_{p}$ Top-1 Test Acc. Train Loss. Network Method $Q_{p}$ Top-1 Test Acc. Train Loss. ResNet-56 [8] AdamW [19] N/A 0.9108 ± 0.0045 0.0057 ± 0.0013 ViT-7/8/8-384 [5] AdamW [19] N/A 0.8398 ± 0.0028 0.0087 ± 0.0092 SAM [6] N/A 0.9160 ± 0.0021 0.0221 ± 0.0051 SAM [6] N/A 0.8432 ± 0.0032 0.0273 ± 0.0101 ZSharp (Ours) 0.95 0.9264 ± 0.0032 0.0630 ± 0.0064 ZSharp (Ours) 0.95 0.8543 ± 0.0029 0.0647 ± 0.0216 ZSharp (Ours) 0.90 0.9212 ± 0.0015 0.0710 ± 0.0061 ZSharp (Ours) 0.90 0.8482 ± 0.0031 0.0748 ± 0.0081 ZSharp (Ours) 0.85 0.9189 ± 0.0023 0.0679 ± 0.0067 ZSharp (Ours) 0.85 0.8424 ± 0.0043 0.0825 ± 0.0086 ZSharp (Ours) 0.80 0.9153 ± 0.0027 0.0731 ± 0.0053 ZSharp (Ours) 0.80 0.8421 ± 0.0038 0.0863 ± 0.0119 ZSharp (Ours) 0.75 0.9132 ± 0.0017 0.0789 ± 0.0079 ZSharp (Ours) 0.75 0.8378 ± 0.0027 0.0999 ± 0.0102
Table 2: Training hyperparameters and results for all experiments, with settings identical to those in the Experimental Settings section except for varying $Q_{p}$ values. Models include ResNet-56 [8] and ViT-7/8/8-384 [5], where ViT-7/8/384 denotes a Vision Transformer with 7 layers, 8 attention heads, a patch size of 8, and an MLP dimension of 384, all trained without pre-trained weights.
Table 2 shows ZSharp achieves the highest Top-1 Test Accuracy at $Q_{p}=0.95$, with 0.9264 ± 0.0032 on ResNet-56 [8] and 0.8543 ± 0.0029 on ViT-7/8/8-384 [5], outperforming AdamW [19] and SAM [6]. As $Q_{p}$ decreases to 0.75, test accuracy nears SAM’s (e.g., 0.9132 on ResNet-56), reflecting ZSharp’s alignment with SAM’s behavior. Based on these results, we identify $Q_{p}=0.95$ as the optimal value and use it for all subsequent experiments to maximize generalization performance.
5 Discussion
Generalization Effect. Figure 5 shows that ZSharp exhibits a higher train loss than other SAM variants (Baseline, SAM, Friendly-SAM, ASAM) for ResNet-56 and ResNet-110 on CIFAR-10, yet achieves superior test accuracy.
Figure 5: Train Loss comparison on CIFAR-10 for ResNet-56 (left) and ResNet-110 (right) across different SAM variants: Baseline, SAM, Friendly-SAM, ASAM, and ZSharp (Ours).
This aligns with Figure 4, where ZSharp’s update path more effectively navigates toward flatter minima, enhancing generalization. The higher train loss indicates reduced overfitting, as ZSharp avoids memorizing training patterns that impair generalization [30]. Instead, Z-score filtering prioritizes significant gradient components, improving generalization to unseen data. Prior studies support this behavior: [2] note that higher train loss with flatter minima can yield robust generalization, while [20] show that models with higher train loss but lower generalization gaps often perform better on test data. Similarly, [12] find that higher train loss from large-batch training can lead to better generalization by finding wider minima. Thus, ZSharp’s higher train loss and superior test accuracy reflect improved generalization performance, highlighting the efficacy of Z-score-based gradient filtering.
Proof of Convergence for Z-Sharp. We observe that ZSharp constitutes a strict generalization of SAM, parameterized by the percentile threshold $Q_{p}$. When $Q_{p}=0$, ZSharp is equivalent to SAM, and the update dynamics follow the same convergence behavior as previously analyzed in sharpness-aware optimization literature. Conversely, when $Q_{p}=1$, the perturbation vanishes, and the update reduces to a standard first-order optimizer (e.g., SGD or Adam), whose convergence guarantees are well established. For intermediate values $10<Q_{p}<$ 1, ZSharp introduces a filtered ascent step while preserving the descent structure of SAM. Although a formal convergence proof for these intermediate cases is nontrivial due to the discontinuous nature of the filtering mask, we posit that the optimization dynamics remain stable under mild conditions. A rigorous theoretical analysis of this regime is left as an important direction for future work.
6 Conclusion
In this paper, we proposed ZSharp, a novel extension of Sharpness-Aware Minimization (SAM) that incorporates statistical filtering via Z-score normalization to enhance generalization in deep neural network training. By selectively retaining only the most statistically significant gradient components during the ascent phase, ZSharp effectively suppresses noisy updates and better aligns perturbations with sharpness-relevant directions in the loss landscape.
Unlike prior SAM variants, ZSharp requires only one additional hyperparameter—the percentile threshold—yet remains computationally lightweight and fully compatible with existing SAM pipelines and optimizers. Extensive experiments across various datasets (CIFAR-10, CIFAR-100, Tiny-ImageNet) and architectures (ResNet, VGG, ViT) consistently demonstrate that ZSharp improves test accuracy while tolerating higher training loss, indicating superior generalization.
Our theoretical formulation shows that ZSharp reduces to SAM under certain conditions, establishing it as a generalization of SAM. These findings support the utility of statistically guided gradient filtering for curvature-aware optimization, especially in high-dimensional or noisy regimes. Future work includes extending ZSharp to large-scale pretraining, robust learning, and federated settings.
7 Acknowledgements
This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.